OpenStudy (anonymous):

The probability of experiencing a significant personnel injury on the shop floor of a manufacturing plant in any given work day is 1%. What is the probability of having zero significant personnel injuries during a period of 30 work days? a) 0.7000 b) 0.6690 c) 0.7397 d) 0.6050 e) 0.9900 What is the probability of having 1 or more significant personnel injuries in that time? a) 0.3310 b) 0.3950 c) 0.9900 d) 0.5052 e) 0.2603

OpenStudy (anonymous):

Binomial formula should work here.

OpenStudy (kropot72):

\[P(zero\ injuries)=\left(\begin{matrix}30 \\ 0\end{matrix}\right)0.01^{0}(0.99)^{30}=?\] \[P(1\ injury)=\left(\begin{matrix}30 \\ 1\end{matrix}\right)0.01^{1}(0.99)^{29}\] \[P(2\ injuries)=\left(\begin{matrix}30 \\ 2\end{matrix}\right)0.01^{2}(0.99)^{28}\] \[P(3\ injuries)=\left(\begin{matrix}30 \\ 3\end{matrix}\right)0.01^{3}(0.99)^{27}\] \[P(4\ injuries)=\left(\begin{matrix}30 \\ 4\end{matrix}\right)0.01^{4}(0.99)^{26}\] To find the probability of 1 or more injuries in 30 days, sum the probabilities of 1, 2, 3 and 4 injuries.